ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 90C20,90C30,90C06,65K05 Nowak, Ivo Bézier-Semidefinite Relaxations of nonconvex quadratic programs over a simplex Some relaxations of nonconvex quadratic programs over a simplex are proposed. The approach is based on quadratic approximations of the convex envelope using Bézier methods and semidefinite programming which results in a convex quadratic program (QP). The optimal value of QP is a lower bound for the original problem. A minimizer of QP is an estimation of a global minimizer of the original problem.